OpenStudy (anonymous):

Find the equation of the tangent line of y=5^x and x=1. Round your answer to one decimal place. I'm guessing I need to know y'=ln(5) * 5^x? But I'm not sure what to do after that.

OpenStudy (anonymous):

So at x = 1, y' = 5ln5

OpenStudy (anonymous):

That means the slope at x = 1 is 5 ln 5

OpenStudy (anonymous):

So the equation of the line at x = 1 (y = 5) is y - 5 = 5ln5(x-1)

OpenStudy (anonymous):

Since this needs to equal to y, would that be... y-5 = 5ln5(x-1) y= 5ln5(x-1)+5 But it's telling me it's wrong.

OpenStudy (anonymous):

Can you just distribute normally? y= 5 ln (5) (x-1) + 5 I'm guessing ln (5) needs to stay in tact...

OpenStudy (anonymous):

yes, ln(5) stays intact as thats a single number called ln(5).

OpenStudy (anonymous):

I just left the equation in point-slope form. You can solve for y and write the equation in slope-intercept form.

OpenStudy (anonymous):

I'm not sure what to do, really. I know in slope-intercept it would be y=1/8x - 3.0 but I don't know how to keep that ln(5) intact... I have one more chance for this problem.

OpenStudy (anonymous):

Use 1.6 for ln(5)....as in the directions of this question, it asks to use one decimal place.

OpenStudy (anonymous):

Oh, okay. I got it. Thanks!